# On uniqueness in structured model learning

### v2

@@ Appendix E Example on choice of regularization
$\|f_{\tilde{\theta}^{m}}\|_{L^{2}(0,1)}^{2}+\|\nabla f_{\tilde{\theta}^{m}}\|_{L^{\infty}(0,1)}+\lambda_{0}m^{3/2}\|f_{\tilde{\theta}^{m}}(u_{m})-g_{m}\|_{L^{2}(0,1)}^{2}+\mu^{m}\|\tilde{\theta}^{m}\|\\ \leq 1+\lambda_{0}(3m^{1/2})^{-1}+\mu^{m}\|\hat{\theta}^{m}\|.$ (62)
As a consequence, there exists a weakly-convergent subsequence of $(f_{\tilde{\theta}^{m}})_{m}$ with limit $\tilde{f}\in W^{1,\infty}(0,1)$ that by the Rellich-Kondrachov embedding converges strongly in $\mathcal{C}(0,1)$, w.l.o.g. for the entire sequence. Employing (weak) lower semicontinuity of the respective norms and (62) imply that
$\|\tilde{f}\|_{L^{2}(0,1)}^{2}+\|\nabla\tilde{f}\|_{L^{\infty}(0,1)}\leq 1=\|\tilde{f}^{\dagger}\|_{L^{2}(0,1)}^{2}+\|\nabla\tilde{f}^{\dagger}\|_{L^{\infty}(0,1)}.$
We show that $\tilde{f}$ fulfills the constraint of ($\mathcal{P}^{\dagger}_{W^{1,\infty}}$), i.e. $\tilde{f}(0)=1$ which by uniqueness of the solution of ($\mathcal{P}^{\dagger}_{W^{1,\infty}}$) implies $\tilde{f}=\tilde{f}^{\dagger}$ as claimed. Indeed
$\|\tilde{f}(0)-1\|_{L^{2}(0,1)}\leq\|\tilde{f}(u)-\tilde{f}(u_{m})\|_{L^{2}(0,1)}+\|\tilde{f}(u_{m})-f_{\tilde{\theta}^{m}}(u_{m})\|_{L^{2}(0,1)}\\ +\|f_{\tilde{\theta}^{m}}(u_{m})-g_{m}\|_{L^{2}(0,1)}+\|g_{m}-1\|_{L^{2}(0,1)}.$ (63)
The fourth term converges to zero by the previous considerations. The third term approaches zero by (62). For the second term note that
$\|\tilde{f}(u_{m})-f_{\tilde{\theta}^{m}}(u_{m})\|_{L^{2}(0,1)}^{2}=\int_{0}^{1}|\tilde{f}(u_{m}(x))-f_{\tilde{\theta}^{m}}(u_{m}(x))|^{2}\mathop{}\!\mathrm{d}x\leq\|\tilde{f}-f_{\tilde{\theta}^{m}}\|_{L^{\infty}(0,1)}^{2}$
which converges to zero as $m$ grows. Finally also the first term converges to zero since
$\|\tilde{f}(u)-\tilde{f}(u_{m})\|_{L^{2}(0,1)}^{2}=\int_{0}^{1}|\tilde{f}(u(x))-\tilde{f}(u_{m}(x))|^{2}\mathop{}\!\mathrm{d}x\leq\|\nabla\tilde{f}\|_{L^{\infty}(0,1)}^{2}\|u-u_{m}\|_{L^{2}(0,1)}^{2}.$
@@
To show that the above counterexample is also observable in practice, we implemented it numerically using single-hidden layer ReLU-activated neural networks with input- and output dimension one and $m\in\mathbb{N}$ nodes in the hidden layer with enabled bias term (to be able to represent constant functions exactly). The network training for ($\mathcal{P}^{m}_{L^{2}}$) and ($\mathcal{P}^{m}_{W^{1,\infty}}$) was performed over $700$ epochs for $m\in\{15,50,200,1000\}$ using the Adam optimizer with learning rate $2\cdot 10^{-4}$ and weight decay $\mu^{m}=m^{-1}$ for $200$ uniformly sampled training points in the unit interval. Furthermore, we chose $\lambda_{0}=10$. The gradient $\nabla f_{\theta}$ in view of the loss calculation of ($\mathcal{P}^{m}_{W^{1,\infty}}$) is approximated via $30$ uniformly sampled approximations of the gradient using finite differences. An illustration of the results is given in Figure 1.
@@
In Figure 1 the first and second line of subplots correspond to the considerations on ($\mathcal{P}^{m}_{L^{2}}$) and ($\mathcal{P}^{m}_{W^{1,\infty}}$), respectively. The first column of subplots depicts approximations of hidden physics after $700$ epochs for the different $m$ above. In the second column the corresponding losses are plotted. Finally, in the third column the $L^{2}$-deviation of the approximated hidden physics from the unique solutions of the limit problem ($\mathcal{P}^{\dagger}_{L^{2}}$) and ($\mathcal{P}^{\dagger}_{W^{1,\infty}}$) is depicted, respectively. The main conclusion is that under $L^{2}$-regularization an incorrect physical term is reconstructed, whereas with the $W^{1,\infty}$-type penalty the hidden physics is identified correctly. For the latter it is further clear that the unknown term is approximated better for increasing $m$ and the parameterized approximations are relatively flat due to the gradient penalty.
@@
$L^{p}(I;X),\mathcal{C}(I;X)$. Bochner space [50, Section 1.5]
$W^{1,p,q}(I;X)$. Sobolev-Bochner space [50, Section 7.1]

### v3

@@ Appendix E Example on choice of regularization
$\|f_{\theta^{m}}\|_{L^{2}(0,1)}^{2}+\|\nabla f_{\theta^{m}}\|_{L^{\infty}(0,1)}\\ \leq\lambda_{0}m^{1/2}\left[\frac{1}{2}-\left(\left(\frac{1}{2}-\frac{2c}{\pi^{2}m^{1/2}}+\frac{c^{2}}{48m}\right)^{1/2}-\frac{c}{\sqrt{32}\pi m^{1/2}}\right)^{2}\right]+\mu^{m}\|\hat{\theta}^{m}\|.$
The right hand side converges to the constant $\frac{c\lambda_{0}}{4\pi}\left(\frac{8}{\pi}+1\right)$ as $m\to\infty$. Thus, there exist $M\in\mathbb{N}$ and a constant $\alpha>0$ such that for all $m\geq M$ it holds true that
$\|\nabla f_{\theta^{m}}\|_{L^{\infty}(0,1)}\leq\alpha.$
Repeating the arguments starting from (67), but now with the refined bound $\|\nabla f_{\theta}\|_{L^{\infty}(0,1)}\leq\alpha$ on the gradients, yields for $m\geq M$ that
$\|f_{\theta^{m}}\|_{L^{2}(0,1)}^{2}+\|\nabla f_{\theta^{m}}\|_{L^{\infty}(0,1)}\\ \leq\lambda_{0}m^{1/2}\left[\frac{1}{2}-\left(\left(\frac{1}{2}-\frac{2\alpha}{\pi^{2}m}+\frac{\alpha^{2}}{48m^{2}}\right)^{1/2}-\frac{\alpha}{\sqrt{32}\pi m}\right)^{2}\right]+\mu^{m}\|\hat{\theta}^{m}\|,$
which converges to zero as $m\to\infty$. As a consequence, it holds true that
$\lim_{m\to\infty}\|\nabla f_{\theta^{m}}\|_{L^{\infty}(0,1)}=0\quad\text{and}\quad\lim_{m\to\infty}\|f_{\theta^{m}}\|_{L^{2}(0,1)}=0,$
implying that $(f_{\theta^{m}})_{m}$ converges uniformly to zero, the unique solution to ($\mathcal{P}^{\dagger}_{W^{1,\infty}}$).
@@
To show that the above counterexample is also observable in practice, we implemented it numerically using two-hidden layers neural networks with input- and output dimension one. The first layer consists of $10$ nodes and is sine-activated with frequency $2$ (to enable the representation of high frequency oscillations). The second layer consists of $m\in\mathbb{N}$ nodes and is ReLU-activated. The network training for ($\mathcal{P}^{m}_{L^{2}}$) and ($\mathcal{P}^{m}_{W^{1,\infty}}$) was performed over $1500$ epochs for $m\in\{10,100,1000\}$ using the Adam optimizer with learning rate $3\cdot 10^{-3}$ and weight decay $\mu^{m}=0.1\cdot m^{-1}$ for $200$ uniformly sampled training points in the unit interval. Furthermore, we chose $\lambda_{0}=1/2$. The gradient $\nabla f_{\theta}$ in view of the loss calculation of ($\mathcal{P}^{m}_{W^{1,\infty}}$) is approximated via $30$ uniformly sampled approximations of the gradient using finite differences. An illustration of the results is given in Figure 1.
@@
In Figure 1 the first and second line of subplots correspond to the considerations on ($\mathcal{P}^{m}_{L^{2}}$) and ($\mathcal{P}^{m}_{W^{1,\infty}}$), respectively. The first column of subplots depicts approximations of hidden physics after $1500$ epochs for the different $m$ above. In the second column the corresponding data term losses are plotted. Finally, in the third column the $L^{2}$-deviation of the approximated hidden physics from the unique solutions of the limit problem ($\mathcal{P}^{\dagger}_{L^{2}}$) and ($\mathcal{P}^{\dagger}_{W^{1,\infty}}$), the zero function, is depicted, respectively. The main conclusion is that under $L^{2}$-regularization the physical term reconstruction fails, whereas with the $W^{1,\infty}$-type penalty the hidden physics is identified correctly. For the latter it is further clear that the unknown term is approximated better for increasing $m$ and the parameterized approximations are relatively flat due to the gradient penalty. Note that the data term loss of ($\mathcal{P}^{m}_{W^{1,\infty}}$) stagnates as predicted by our analysis above.
@@
$L^{p}(I;X),\mathcal{C}(I;X)$. Bochner space [98, Section 1.5]
$W^{1,p,q}(I;X)$. Sobolev-Bochner space [98, Section 7.1]